OpenStudy (shaleiah):

A super-deadly strain of bacteria is causing the zombie population to double every 2 days. Currently, there are 25 zombies. After how many days will there be 25,600 zombies?

psirockin2 (psirockin2):

Okay so... it's increasing by double every 2 days

psirockin2 (psirockin2):

8 days

OpenStudy (shaleiah):

how did you get 8 days?

OpenStudy (shaleiah):

@psirockin2

OpenStudy (amorfide):

25, 50, 100, 200, 400,...., 25600 you have a common ratio of 50/25=100/50=200/100 etc so r=2 geometric sequence is a, ar, ar^2, ar^3 so you want ar^n=25600 our first term is 25, r=2, so \[25(2^{n})=25600\]

OpenStudy (princeharryyy):

that's a GP. Geometric progression.

OpenStudy (amorfide):

please not that for a geometric sequence a,ar,ar^2,....,ar^n is the standard form for it a is the first term of the sequence r is the ratio to get the ratio, r, you would divide 2 consecutive terms ar/a=ar^2/ar=ar^3/ar^2, etc sorry I thought I would clarify just incase

OpenStudy (princeharryyy):

@amorfide don't u thing u might still be wrong.

OpenStudy (princeharryyy):

day 0 no. of bacteria = 25 after 2 days bacteria's = 50 after another 2 days bacteria's = 100 after another 2 days bacteria's = 200 after another 2 days bacteria's =400 after another 2 days bacteria's =800 after another 2 days bacteria's =1600 after another 2 days bacteria's =3200 after another 2 days bacteria's =6400 after another 2 days bacteria's =12800 after another 2 days bacteria's 28600 total days add all the days. 2*10 = 20

OpenStudy (princeharryyy):

your aren't wrong, that's a GP only.

OpenStudy (amorfide):

ah, yeah, sorry

OpenStudy (princeharryyy):

@amorfide

OpenStudy (princeharryyy):

the only problem with that is it's not after every day that's after every two days.

OpenStudy (princeharryyy):

so, u wud have to keep in mind of that while solving that. otherwise it's perfect np.

OpenStudy (amorfide):

so since I treated the progression as per day, I should multiply the answer by 2, so no problem. Thank you for correcting

OpenStudy (princeharryyy):

yeah. that's what I am saying :)

OpenStudy (amorfide):

I will rewrite it for the op

OpenStudy (shaleiah):

I'm looking for the growth rate formula

OpenStudy (amorfide):

no idea what that is

OpenStudy (princeharryyy):

@amorfide gave u the formula. but u wud have to keep in mind of the days.

OpenStudy (princeharryyy):

that's rate only.

OpenStudy (amorfide):

I can't help @princeharryyy please would you be so kind

OpenStudy (shaleiah):

@563blackghost